# Self-organizing railway traffic management

### v1

@@ I Introduction
Railway traffic is often perturbed by unexpected events, which require train routes and schedules to be carefully selected to limit delay propagation. This selection is currently made by dispatchers, with very limited decision support tools. From the traffic control centers, they make decisions on traffic evolution in a control area based on experience and a set of simple heuristic rules. The railway community, including public authorities, industry and academia, is today very actively aiming at the introduction of intelligent decision support tools in traffic control centers. Among other initiatives, the European Joint Undertakings Shift2Rail [32] and EuropeRail [11] are remarkable examples of this shared aim. In virtually all considered possibilities for such an introduction, the structure of the decision-making process is not questioned: optimization [5, 3, 12, 20], and recently, machine learning techniques [19, 16, 41] are proposed to support dispatchers in their centralized decisions.
@@
A major change of paradigm in this process has recently made the object of a few studies, i.e., a Self-Organizing Traffic Management System (SO-TMS) in which trains autonomously make decisions on their routes and schedules to deal with perturbations [22]. Indeed, self-organizing systems have shown to be effective in many contexts, achieving the most diverse objectives. In the railway traffic management context, self-organization may bring a number of benefits. First, thanks to the ability to make local decisions that result in good overall system performance, the system would quickly react to perturbations, thus shortening recovery times. The local decision-making would also allow an efficient scaling up to large networks, while most TMS approaches can hardly do so [39]. Second, an SO-TMS may strongly increase the system flexibility, as trains can base their decisions on factors as real-time multimodal traffic conditions or demand flows, possibly disregarding fixed timetables. Third, it would satisfy the needs of a railway market in which different railway companies compete for the use of the same tracks, or even for the provision of equivalent services. In this case, it may not be acceptable to share with dispatchers sensitive information determining delay costs, such as the number of passengers onboard or the value of the carried freight. Hence, in a classic centralized system, similar costs are never considered, limiting decision-making to minimization of delays. In a self-organizing system, instead, these known costs do not need to be shared, and trains may negotiate traffic evolutions based on their actual interests. Despite these possible benefits, to the best of our knowledge, no such system has ever been convincingly applied to railway traffic management so far. Initial studies on TMS based on train decision-making have been proposed. A first approach proposes that trains decide their movements in a loop-shaped railway line, based on the behavior of the train ahead [34]. Another approach instructs trains to decide on their passing order at a junction based on swarm intelligence principles: trains are grouped in platoons and make decisions as a swarm, by applying rather simple rules after checking for the infrastructure occupation status [40]. A third approach focuses on resolving conflicts individually by adjusting train schedules or routes [38].
@@
This paper builds upon the framework discussed above [7] by providing concrete designs and implementations. Specifically, we introduce a spatio-temporal perspective for neighborhood selection, a variant of the state-of-the-art centralized algorithm RECIFE-MILP [29] for hypothesis generation, a path-based hypothesis compatibility verification with hypothesis sharing, a voter-model-inspired consensus-seeking approach, and a novel RECIFE-MILP variant for merging selected hypotheses into the global plan. To the best of our knowledge, this is the first proposal of a complete system based on traffic self-organization capable of consistently producing feasible and optimized traffic management plans in complex networks and quickly reacting to perturbations.
We thoroughly evaluate our approach on a case study representing traffic in a busy portion of the Italian network, between the cities of Segrate and Ospitaletto in Lombardy, Italy. Here, mixed traffic operated by different railway companies share the available capacity, posing significant challenges related to competition, fairness, and the efficient utilization of limited infrastructure. Our results demonstrate that, in these instances, the proposed self-organizing approach consistently outperforms the state-of-the-art centralized algorithm RECIFE-MILP [29]. Specifically, we compare the two approaches in terms of total delay suffered by trains in a realistic deployment, exploiting a microscopic railway simulator to reproduce traffic. In summary, the main contribution of this paper is the proposal of the first comprehensive self-organization approach for railway traffic management. In addition to the formalization and implementation of the algorithms composing this approach, we also propose an articulated experimental analysis to assess its applicability and performance in a possible future deployment.
@@ III Real-time Railway Traffic Management
In this paper, we consider a deployment in which a closed loop is enacted between the field and the decision making process [31, 4, 38], be it self-organized or centralized. Here, the traffic evolution is controlled through the Real-Time Traffic Plan (RTTP) [31], which indicates train routes and passing orders. This evolution is constantly monitored and, periodically, reported in the form of a Traffic State (TS) [31]. The decision making process considers the TS and produces a new RTTP to react to the possibly observed perturbations.
@@ IV Self-organization process
To self-organize, trains need to identify the traffic share with which they may interact in the near future, then to generate hypotheses on how these shares may evolve, to finally assess other trains’ hypotheses and get to an agreement on which ones shall be implemented. This process has been formally defined in [7] through the definition of five modules (see Figure 1).
@@
When trains receive the current RTTP and TS, they independently carry out a neighborhood identification to determine those trains they could be in conflict with in the near future, and that therefore need to be taken into account for traffic management. Then, each train executes the hypothesis generation to formulate possible traffic management strategies involving all its neighbors. Subsequently, it shares these hypotheses with its neighbors and carries out an hypothesis compatibility check to verify which of the hypotheses received from its neighbors could be selected concurrently with its own without generating feasibility issues. In the last step carried out independently by each train—the hypotheses selection—each train engages with its neighbors in an iterative process, during which they all try to reach an agreement on the hypotheses to select, minimizing their own costs while preserving overall compatibility with neighbors. Finally, after the hypothesis selection, the trains share their selected hypotheses with the traffic control center, which is in charge of the merge of all hypotheses into an RTTP. If incompatibilities are detected at this stage, which may specifically happen for trains that are currently too far to be included in each other neighborhoods, the control center repairs the RTTP by resolving the detected conflicts without modifying the selected hypotheses in the short term.
@@
Figure 2 depicts an example of application of this process. Here, four trains travel in a control area. After neighborhood identification, it emerges that A and B have the same neighborhood (A, B and C), C includes all trains in its neighborhood (A, B, C, D) while D only has C in its neighborhood (D, C). Then, each train generates traffic management hypotheses and sends them to its neighbors for the compatibility check. Subsequently, a hypothesis selection process takes place whereby each train iteratively chooses the preferred hypothesis depending on its corresponding cost and its compatibility with the hypotheses selected by its neighbors. After this selection, the chosen hypotheses are sent to the traffic control center and merged in the new RTTP.
@@ V-B Hypothesis generation
To generate different hypotheses of traffic management strategies, each train $t$ considers its own neighborhood and tries to optimize schedules and routes to minimize delays. To this end, a variant of the RECIFE-MILP algorithm for the rtRTMP [29] is employed, first proposed in [7]. In particular, the algorithm is based on the solution of a mixed-integer linear programming (MILP) formulation and it operates in two phases. First, train schedules are optimized only considering the routes of the input RTTP. Second, this solution is used as a warm start and all route alternatives are considered. In the considered variant, routes and schedules are optimized only for the focal train $t$ and its neighbors $t^{\prime}\in\mathcal{N}(t)$. The paths of all the other trains are considered as hard constraints: the routes must remain the ones of the RTTP, and the schedules may be changed provided that the trains exit the control area no later than what expected before the optimization. The algorithm stops as soon as a solution is proven optimal or when the available computational time has elapsed.
The objective function minimized is the weighted sum of train delays at their exit from the control area. The weights represent the cost of delays for each train. They are input to the algorithm and they depend on the context of interest. For example, if a single company operates trains in a control area, the weights would depend on the company’s policy: all trains could be equal, or some of them may be perceived as more important. Also, the weights would be known by all trains, as they would share the same objective function. Instead, if competing companies operate, trains might set these weights based on their own cost of delay, and considering an educated guess for the other trains. Whatever the context, we assume that all trains are interested in the good functioning of the whole system, in addition to limiting their own delay. However, let us remark that any linear function of delay may be used in different scenarios.
The considered RECIFE-MILP variant produces a number of different solutions. Each solution computed by train $t$ is a hypothesis $h_{t}$. Then, at most $\bar{h}$ solutions are retained, those that are not further than $p\%$ from the optimal one or the best one identified by the algorithm, with $p$ and $\bar{h}$ input parameters. We refer to the set of hypotheses generated for train $t$ as $H_{t}$.
@@ V-C Hypothesis compatibility
A pair of hypotheses $h_{1}$ and $h_{2}$ generated by two neighboring trains $t_{1}$ and $t_{2}$ may differ in multiple aspects related to scheduling and routing. However, despite differences, they may contain plans for $t_{1}$ and $t_{2}$ that, if implemented, produce no infeasibility issues. In this sense, $h_{1}$ and $h_{2}$ are said to be compatible. Specifically, to perform the compatibility check, the route of the train $t_{1}$ from $h_{1}$ and the one of train $t_{2}$ from $h_{2}$ are confronted. If there exists at least one TDS in which utilization times overlap among $t_{1}$ and $t_{2}$, then $h_{1}$ and $h_{2}$ are declared incompatible.
@@ V-D Hypothesis selection
The hyperparameter $K$ tunes a speed-accuracy trade off. High values of $K$ imply that the train considers multiple (possibly all) neighbors during the decision making process. This can lead a train to simultaneously seek compatibility with a large set of neighbors, hence increasing the speed of convergence towards a shared solution, notwithstanding the cost of the compatible hypotheses. Conversely, when $K$ is small (and possibly, $K=1$), the train considers only a small set of neighbors at a time, and therefore the speed of convergence may be slower. However, in these conditions, a train has a higher chance of selecting among multiple compatible hypotheses, hence preferring those with lower cost. To maximize decision accuracy and speed, we propose an adaptive strategy that adjusts the value of the parameter $K$ over time. Trains start the consensus process with the highest possible value of $K$ (i.e., $K$ equal to the total number of neighbors) and gradually decrease it towards $K=1$ from iteration to iteration. In this way, the self-organized process allows to rapidly move towards agreement with neighbors, and to gradually improve decision quality once trains start being able to prioritize hypotheses with smaller cost. In [8], we extensively investigated the role of $K$ and compared various strategies, concluding that the adaptive approach features both the highest converge speed and the highest rate of convergence to an optimal solution.
@@ V-E Merge
Once this merging is performed, a consistency check is carried out on the updated RTTP to identify TDSs where paths overlap, if any. This may happen if two trains that are too far to be in the same neighborhood plan to travel across the same TDS at a later stage. In other words, compatible hypotheses only ensure that conflicts do not exist in the short-term, but overlaps are possible in the long run. If this happens, a repair procedure is executed exploiting a variant of the RECIFE-MILP algorithm [29], in which train routes are preserved, as well as train schedules for which a consensus was found. In particular, this is the case for all scheduling decisions that are realized within the time horizon $T_{h}$ of the neighborhood selection. All other scheduling decisions are optimized, considering the minimization of the unweighted total delay. In fact, the control center has no access to the weights assigned by the railway companies to their trains.
Remark that the performance of the merging phase at the control center diverges from typical, strictly self-organizing processes, where all decisions are based solely on local information. However, a global check is the only option to guarantee the overall feasibility of the RTTP. As we are proposing here an approach which may in principle be deployed in practice, supplying an overall feasibility guarantee is a necessary condition. It could be performed by a single train, but we posit that the traffic control center is a more sensible choice to ensure the fairness of the repair operation, if necessary, and remain to some extent consistent with the current traffic management organization.

### v2

@@ I Introduction
Railway traffic is often affected by unexpected events. These events can be perturbations or disturbances, i.e., minor deviations from the timetable, or disruptions, i.e., major degrading events [6, 30]. While disruptions impose significant changes to the timetable to be agreed upon by different stakeholders, perturbations require train routes and schedules to be carefully selected to limit delay propagation. This selection is currently made by dispatchers, with very limited decision support tools. To deal with perturbations, from the traffic control centers, dispatchers make decisions on traffic evolution in a control area based on experience and a set of simple heuristic rules. The railway community, including public authorities, industry and academia, is today very actively aiming at the introduction of intelligent decision support tools in traffic control centers. Among other initiatives, the European Joint Undertakings Shift2Rail [42] and EuropeRail [18] are remarkable examples of this shared aim. In virtually all considered possibilities for such an introduction, the structure of the decision-making process is not questioned: optimization [6, 11, 19, 27], and recently, machine learning techniques [26, 23, 51] are proposed to support dispatchers in their centralized decisions.
@@
A major change of paradigm in this process has recently made the object of a few studies, i.e., a Self-Organizing Traffic Management System (SO-TMS) in which trains autonomously make decisions on their routes and schedules to deal with perturbations [29]. Indeed, self-organizing systems have shown to be effective in many contexts, achieving the most diverse objectives. In the railway traffic management context, self-organization may bring a number of benefits. First, thanks to the ability to make local decisions that result in good overall system performance, the system would quickly react to perturbations, thus shortening recovery times. The local decision-making would also allow an efficient scaling up to large networks, while most TMS approaches can hardly do so [49]. Second, an SO-TMS may strongly increase the system flexibility, as trains can base their decisions on factors as real-time multimodal traffic conditions or demand flows, possibly disregarding fixed timetables.
Third, it would satisfy the needs and logic of a competitive railway market, as the one envisaged in Europe [1, 3, 16]. In such a market, railway companies compete for the use of the same tracks, which are controlled by infrastructure managers, typically one per country. In this competitive market, companies may not be willing to share information on the value they attribute to specific trains and on the cost of their delay, since it would disclose sensitive data, such as passenger volumes [8], to competitors. Hence, in a classic centralized system, these costs are unknown by dispatchers and cannot be considered. This lack of knowledge limits decision-making to the plain minimization of delays.
In a self-organizing system, instead, these known costs do not need to be shared, and trains may negotiate traffic evolutions based on their actual interests. Despite these possible benefits, to the best of our knowledge, no such system has ever been convincingly applied to railway traffic management so far. Initial studies on TMS based on train decision-making have been proposed. A first approach proposes that trains decide their movements in a loop-shaped railway line, based on the behavior of the train ahead [44]. Another approach instructs trains to decide on their passing order at a junction based on swarm intelligence principles: trains are grouped in platoons and make decisions as a swarm, by applying rather simple rules after checking for the infrastructure occupation status [50]. A third approach focuses on resolving conflicts individually by adjusting train schedules or routes [48].
@@
This paper builds upon the framework discussed above [13] by providing concrete designs and implementations. Specifically, we introduce a spatio-temporal perspective for neighborhood selection, a variant of the state-of-the-art centralized algorithm RECIFE-MILP [38] for hypothesis generation, a path-based hypothesis compatibility verification with hypothesis sharing, a voter-model-inspired consensus-seeking approach, and a novel RECIFE-MILP variant for merging selected hypotheses into the global plan. To the best of our knowledge, this is the first proposal of a complete system based on traffic self-organization that is capable of quickly reacting to perturbations by consistently producing feasible and optimized traffic management plans in complex networks.
We thoroughly evaluate our approach on a case study representing traffic in a busy portion of the Italian network, between the cities of Segrate and Ospitaletto in Lombardy, Italy. Here, mixed traffic operated by different railway companies share the available capacity, posing significant challenges related to competition, fairness, and the efficient utilization of limited infrastructure. Our results demonstrate that, in these instances, the proposed self-organizing approach achieves results that are very close to optimality. Moreover, it consistently outperforms the state-of-the-art centralized algorithm RECIFE-MILP [38], even when its performance are boosted by the smart reduction of alternative train routes [37]. Specifically, we compare the approaches in terms of total delay suffered by trains in a realistic deployment, exploiting a microscopic railway simulator to reproduce traffic. In summary, the main contribution of this paper is the proposal of the first comprehensive self-organization approach for railway traffic management. In addition to the formalization and implementation of the algorithms composing this approach, we also propose an articulated experimental analysis to assess its applicability and performance in a possible future deployment.
@@ II Self-organization and traffic management
In summary, the works that propose traffic management approaches in which trains are seen as individual decision makers mostly do so as a way to decompose the overall problem [26, 50, 48, 25, 32]. Here, all trains aim at the objective of minimizing the same function of total delay. When one of those approaches allows, at least in principle, different objectives to be considered by different trains [44, 13], the corresponding paper either focuses only on a specific part of the traffic management process or reports performance assessments based on very simplified traffic representations. These can consist of a basic network [44], a minimal set of trains [50], or a simplified model of train operations [26, 50, 48, 13, 32]. Moreover, apart from the paper which we extend [13], only another study considers train decisions that may be critical to traffic feasibility while also ensuring that feasibility is always preserved [50].
@@ III Real-time Railway Traffic Management
In this paper, we consider a deployment in which a closed loop is enacted between the field and the decision making process [48, 40, 10], be it self-organized or centralized. Here, the traffic evolution is controlled through the Real-Time Traffic Plan (RTTP) [40], which indicates train routes and passing orders. This evolution is constantly monitored and, periodically, reported in the form of a Traffic State (TS) [40]. When a TS is available, the decision making process is executed and produces a new RTTP to react to the possibly observed perturbations.
@@ IV Self-organization process
To self-organize, trains need to identify the traffic share with which they may interact in the near future, then to generate hypotheses on how these shares may evolve, to finally assess other trains’ hypotheses and get to an agreement on which ones shall be implemented. This process has been formally defined in [13] through five modules (see Figure 1).
@@
When trains receive the current RTTP and TS, they independently carry out a neighborhood identification to determine the trains they could be in conflict with in the near future, and that therefore need to be taken into account for traffic management. Then, each train executes the hypothesis generation to formulate possible traffic management strategies involving all its neighbors. Subsequently, it shares these hypotheses with its neighbors and carries out a hypothesis compatibility check to verify which of the hypotheses received from its neighbors could be selected concurrently with its own without generating feasibility issues. In the last step carried out independently by each train—the hypothesis selection—each train engages with its neighbors in an iterative process, during which they all try to reach an agreement on the hypotheses to select, minimizing their own costs while preserving overall compatibility with neighbors. Finally, the trains share their selected hypotheses with the traffic control center, which is in charge of the merge of all hypotheses into an RTTP. If incompatibilities are detected at this stage, which may specifically happen for trains that are currently too far to be included in each other neighborhoods, the control center repairs the RTTP by resolving the detected conflicts without modifying the selected hypotheses in the short term.
@@
Figure 2 depicts an example of its application. Here, four trains travel in a control area. After neighborhood identification, A and B have the same neighborhood (A, B and C), C includes all trains in its neighborhood (A, B, C, D) while D only has C in its neighborhood (D, C). Then, each train generates traffic management hypotheses and sends them to its neighbors for the compatibility check. Subsequently, a hypothesis selection process takes place whereby each train iteratively chooses the preferred hypothesis depending on its corresponding cost and its compatibility with the hypotheses selected by its neighbors. After this selection, the chosen hypotheses are sent to the traffic control center and merged in the new RTTP.
@@ V-B Hypothesis generation
To generate different hypotheses of traffic management strategies, each train $t$ considers its own neighborhood and tries to optimize schedules and routes to minimize delays. To this end, a variant of the RECIFE-MILP algorithm for the rtRTMP [38] is employed, first proposed in [13]. In particular, this algorithm is based on the solution of a mixed-integer linear programming (MILP) formulation and it operates in two phases. First, train schedules are optimized only considering the routes of the input RTTP. Second, this solution is used as a warm start and all route alternatives are considered. In the employed variant, routes and schedules are optimized only for the focal train $t$ and its neighbors $t^{\prime}\in\mathcal{N}(t)$. The paths of all the other trains are considered as hard constraints: the routes must remain the ones of the input RTTP, and the schedules may be changed provided that the trains exit the control area no later than what expected before the optimization. The algorithm stops as soon as a solution is proven optimal or when the available computational time has elapsed.
The objective function minimized is the weighted sum of train delays at their exit from the control area. The weights represent the cost of delays of the focal train for itself and all the other trains. By considering different values for these weights, which are an input of the algorithm, this objective function can model individual train behaviors in different contexts. In particular, it can model extreme competition or extreme cooperation, or any combination of them [41]. For example, each train may be interested in minimizing only its own delay, and it may set its own weight to a positive value, while the weights for all the other trains are zero. Alternatively, each train may value its own delay the most, but it may also wish for a well-performing railway system as a whole. Hence, it may set its own weight ten, for example, times higher than the one for the other trains, but keep all weights strictly greater than zero. Moreover, this objective function can capture the lack of information a train may have on the cost of delay for its competitors [8]. In this case, the weight may represent an educated guess on the other trains’ cost rather than a true value. This is the rationale we adopt to define the costs in the computational analysis (Section VI).
However, let us remark that any linear function of delay may be used in different scenarios.
The considered RECIFE-MILP variant produces a number of different solutions. Each solution computed by train $t$ is a hypothesis $h_{t}$. It includes a path for each train, where a path describes the train utilization of all the TDSs along its route: it can be either the original or a new path for the focal train and the trains in its neighborhood, while it is necessarily the path of the input RTTP, or an extremely similar one, for all the other trains. All these paths describe a complete RTTP, and, as such, all train routes and passing orders. At most $\bar{h}$ solutions are retained, those that are not further than $p\%$ from the optimal one or the best one identified by the algorithm, with $p$ and $\bar{h}$ input parameters. We refer to the set of hypotheses generated for train $t$ as $H_{t}$.
@@ V-C Hypothesis compatibility
A pair of hypotheses $h_{1}$ and $h_{2}$ generated by two neighboring trains $t_{1}$ and $t_{2}$ may differ in multiple aspects related to scheduling and routing. However, despite differences, they may contain paths for $t_{1}$ and $t_{2}$ that, if implemented, produce no infeasibility issues. In this sense, $h_{1}$ and $h_{2}$ are said to be compatible. Specifically, to perform the compatibility check, the paths of $t_{1}$ from $h_{1}$ and the one of $t_{2}$ from $h_{2}$ are confronted. If there exists at least one TDS in which utilization times overlap among $t_{1}$ and $t_{2}$, then $h_{1}$ and $h_{2}$ are declared incompatible.
@@ V-D Hypothesis selection
As discussed below, hyperparameter $K$ controls the trade-off between trying to find a feasible solution quickly and aiming at a high-quality solution from each train’s perspective.
Train $t$ then decides whether to keep its current hypothesis or switch to a more compatible one according to the following policy:
@@
As anticipated above, the hyperparameter $K$ tunes a speed-accuracy trade off. High values of $K$ imply that the train considers multiple (possibly all) neighbors during the decision making process. This can lead a train to simultaneously seek compatibility with a large set of neighbors, hence increasing the speed of convergence towards a shared solution. However, in such conditions, it may be difficult to find a compatible hypothesis, and costs may increase. Conversely, when $K$ is small (and possibly, $K=1$), the train considers only a small set of neighbors at a time, and therefore the speed of convergence may be slower. However, in these conditions, a train has a higher chance of selecting among multiple compatible hypotheses, and convergence to low-cost hypotheses can gracefully build up within the network. To maximize decision accuracy and speed, we propose an adaptive strategy that adjusts the value of the hyperparameter $K$ over time. Trains start the consensus process with the highest possible value of $K$ (i.e., $K$ equal to the total number of neighbors) and gradually decrease it towards $K=1$ from iteration to iteration. In this way, the self-organized process allows to rapidly move towards agreement with neighbors, and to gradually improve decision quality once trains start being able to prioritize hypotheses with smaller cost. In [14], we extensively investigated the role of $K$ and compared various strategies, concluding that the adaptive approach features both the highest convergence speed and the highest rate of convergence to an optimal solution.
Algorithm 1 summarizes the hypothesis selection procedure from the perspective of a focal train.
Algorithm 1 Hypothesis selection procedure of each focal train $t$
1:Input: Hypothesis graph $G_{h}$.
2:Returns: Each train’s selected hypothesis.
[…]
4:while $\exists$ neighor with selected hypotheses incompatible with $h_{t}$
5:    & available time not elapsed do
6:  $t.\mathit{updateNeighborSelectedHypotheses}()$
7:  ${\tilde{\mathcal{N}}(t)}\leftarrow t.\mathit{selectNeighborSubset}(\mathcal{N}(t))$
8:  $t.\mathit{rankHypotheses}(H_{t},{\tilde{\mathcal{N}}(t)})$
9:  if $\chi(h_{t})<K$ then
10:   $h_{t}\leftarrow t.\mathit{selectNewHypothesis}(\boldsymbol{\chi})$
11:  end if
12:  $t.\mathit{shareHypothesisWithNeighbors}()$
13:end while
14:return $h_{t}$
@@ V-E Merge
Once this merging is performed, a consistency check is carried out on the updated RTTP to identify TDSs where paths overlap, if any.
Indeed, when hypotheses are generated, each focal train $t$ considers the paths of trains not belonging to its neighborhood (e.g., $\bar{t}$) as constraints. If these paths are not changed by the respective trains by means of their hypothesis selection, the new path of the focal train $t$ will not raise any inconsistency. However, if $\bar{t}$, as a focal train itself, achieves a consensus with its own neighbors and changes its path, there is no guarantee that the new choice of $t$ and $\bar{t}$ will be consistent. If two or more paths overlap,
a repair procedure is executed exploiting a variant of the RECIFE-MILP algorithm [38], in which train routes are preserved, as well as train schedules for which a consensus was found. In particular, this is the case for all scheduling decisions that are realized within the time horizon $T_{h}$ of the neighborhood selection. All other scheduling decisions are optimized, considering the minimization of the unweighted total delay. In fact, the control center has no access to the weights assigned by the railway companies to their trains.
Remark that the fulfillment of the merging phase at the control center departs from typical, strictly self-organizing processes, where all decisions are based solely on local information. However, a global check is the only option to guarantee the overall feasibility of the RTTP. As we are proposing here an approach which may in principle be deployed in practice, supplying an overall feasibility guarantee is a necessary condition. This global check could be performed by a single train, but we posit that the traffic control center is a more sensible choice to ensure the fairness of the repair operation, if necessary, and remain to some extent consistent with the current traffic management organization.